A rectangular lot whose perimeter is 310 ft is fenced along three sides. an expensive fencing along the lot's length costs $25 per foot. An inexpensive fencing along the two side widths costs only $5 per foot. The total cost of the fencing along three sides comes to $3050. What are the lot's dimensions
Answer by lwsshak3(11628)

A rectangular lot whose perimeter is 310 ft is fenced along three sides. an expensive fencing along the lot's length costs $25 per foot. An inexpensive fencing along the two side widths costs only $5 per foot. The total cost of the fencing along three sides comes to $3050. What are the lot's dimensions.
**
let x=length of lot
let y=width of lot
2x+2y=310
x+y=155
y=155-x
..
25x+5*2y=3050
25x+10(155-x)=3050
25x+1550-10x=3050
15x=1500x
x=100
y=155-x=55
..
ans:
length of lot=100 ft
width of lot=55 ft